# Detecting Spatial Dependence in Transcriptomics Data using Vectorised Persistence Diagrams

### v1

@@ Abstract
Evaluating spatial patterns in data is an integral task across various domains, including geostatistics, astronomy, and spatial tissue biology. The analysis of transcriptomics data in particular relies on methods for detecting spatially-dependent features that exhibit significant spatial patterns for both explanatory analysis and feature selection. However, given the complex and high-dimensional nature of these data, there is a need for robust, stable, and reliable descriptors of spatial dependence. We leverage the stability and multi-scale properties of persistent homology to address this task. To this end, we introduce a novel framework using functional topological summaries, such as Betti curves and persistence landscapes, for identifying and describing non-random patterns in spatial data. In particular, we propose a non-parametric one-sample permutation test for spatial dependence and investigate its utility across both simulated and real spatial omics data. Our vectorised approach outperforms baseline methods at accurately detecting spatial dependence. Further, we find that our method is more robust to outliers than alternative tests using ‘Moran’s I.’
@@ 1 Introduction
Omics data, that is measurements of molecules within an organism or a cell, provide crucial insights into biological processes and cellular function. Specifically, transcriptomics plays a vital role in studying RNA transcripts i.e. the expression of genes via the transcriptome and is thus capable of providing snapshots of cellular function and the mechanisms behind health and disease. Recent advancements in spatial transcriptomics further enable the large-scale analysis of cells’ gene expression and their spatial locations across tissue samples. Leveraging these data, the detection of spatially variable genes (SVGs) is an important analysis step both for explanatory analysis and further downstream tasks, such as spatial domain identification and gene enrichment analysis, uncovering the spatial components of tissue biology [27]. Nevertheless, accurately identifying the spatial patterns in omics data on both local and global scales remains challenging, particularly due to the high-dimensional, sparse and noisy nature of these data. The difficulty of modelling technical and systematic errors, the ongoing debate on even basic preprocessing stpdf, as well as an ever growing understanding of the fundamental processes in cellular biology, therefore necessitate the development of robust and expressive analysis methods [32]. In particular, SVG detection methods should be robust to technical noise and sparsity, correctly identify spatially dependent genes, be independent of the sum of expression values, and show low false positive error rates [13, 24]. In practice, reliable SVG detection methods then help to identify the spatial patterns of disease e.g. by detecting spatial clusters of cancer cells or necrotizing tissue within an organ as well as identifying the genes driving these spatial changes [27].
Meanwhile, topological data analysis (TDA) has given scholars a new toolbox for quantifying complex structures in biological networks at multiple scales motivated by the rigorous mathematical study of shapes [17]. Indeed, when adjusting for noise, molecular data in particular is equipped with meaningful coarse geometry [32]. Capturing both local and global characteristics of the shape of data, TDA is thus poised to quantify spatial patterns in molecular data. Computing persistent homology on spatial graphs gives stable and flexible summaries of the distribution of spatial features [11] summarising both local and global patterns in gene expression. In particular, our work demonstrates how toplogical descriptors can be used to distinguish spatial signals in transcriptomics data and identify genes whose expression varies across space. Our contributions are as follows:
@@ 2 Related Work
2.1 Spatially-Variable Gene Detection
@@
The field of spatial omics data analysis has grown rapidly in the last 10 years fueled by advancements in sequencing technologies and data availability [26]. The term spatial omics denotes spatially-resolved molecular measurements in general, including features extracted from the genome, transcriptome, or proteome. These features are greatly relevant as they give insights into fundamental biological processes. As a key mechanism in biology, the set of DNA encoded in the genome is read out or transcribed by RNA molecules, which make up the transcriptome. Some of these transcripts later code for proteins, which are of great practical importance as they perform the actual biological functions in cells. Omics measurements then give key insights into these fundamental building blocks of biology. Depending on the sequencing method, observations from spatial datasets either correspond to individual cells, enabling the advent of single cell omics, aggregations of multiple cells or even sub-cellular measurements. Associated spatial locations are given by unstructured i.e. continuous coordinates on a tissue sample or spots on a pre-defined grid. Most frequently, features are transcriptomic measurements that summarise the number of times a gene as measured by the number of RNA fragments detected at each capture location [27], the focus of our investigation.
@@
Generally, SVG detection methods fall into numerous categories. Moran’s I is the most established statistical measure for spatial auto-correlation across spatial statistics [25]. As well as its cousin Geary’s C [18], Moran’s I is frequently used as a versatile descriptor of spatial patterns in gene expression values [20, 24]. Despite its simplicity as a mean summary based on weighted variance and covariance estimates, and its well-established theoretical benefits, Moran’s I is known to be sensitive to outliers and noise [40, 10], a disadvantage given the noisy nature of spatial omics data [13]. Further, based on the rich tradition of spatial statistics, there also exists a variety of statistical methods that fit spatial regression models to asses the importance of spatial covariance on gene expression. The most popular of these approaches include SpatialDE and SpatialDE2, which fit Gaussian regression processes, and SPARK or SPARK-X fitting mixed and non-parametric models respectively [14]. However, these models rely on fixed assumptions on the nature of the distribution of gene expression values, which remains a contested question in practice [42]. Lastly, there exist non-parametric methods, such as trendsceek, which models gene expression as a marked point process and uses permutation testing to make inferences [16]. This approach benefits by allowing statistical reasoning under less stringent assumptions but is associated with increased computational costs. Further, model-free frameworks such as sepal, which measures diffusion times, or GLISS, which models graph Laplacians, have also been proposed [1]. However, despite the appeal of such geometric and model-free approaches, multi-scale descriptors, such as persistent homology have not yet been employed in this context. Motivated by the use of geometric descriptors of spatial graphs, we thus aim to further develop persistent homology as a framework for spatial significance testing for spatial omics data analysis.
@@ 2.2 Persistent Homology
Originating from algebraic topology, persistent homology (PH) enables practitioners to measure topological features from data [15] thus describing the shape of data [5]. This is of particular importance for molecular data, which after accounting for noise can be surmised to contain meaningful coarse geometry [32]. PH then gives one such a tool for encoding this geometry in a robust manner.
From a signal processing perspective, persistent homology has demonstrated its utility for detecting peaks in noisy signals guaranteeing stability properties and robustness to small perturbations [22, 21, 17]. The multi-scale nature of filtrations allows practitioners to not only summarise local but also global patterns in spatial data. Persistence-based methods can thus be applied to detect and encode both local and global extrema alike [21]. Indeed, the use of persistent homology as spatial descriptor has been extended to investigating time-varying data and spatial clustering [21, 29]. Thus, persistence has shown its utility for describing spatial patterns, such as local maxima and local minima, in various contexts. It is then only natural to consider the promising potential of using persistence for SVG detection in particular.
From a statistical perspective, the increasingly widespread use of PH for data analysis and budding applications to ML have also been driven by advancements in the statistical theory underpinning the theoretical foundation of PH. In particular, summaries of PH allow for the definition of parametric and non-parametric significance tests and uncertainty estimation. Amongst these approaches, non-parametric techniques, specifically permutation testing, are by far the most utilised and assumption-free methods used for statistical inferences [9, 35]. Arguably, one-sample testing remains somewhat under-explored [11] with most common statistical approaches focusing on two-sample comparisons.
Nevertheless, recent work by Byers et al. [11] demonstrate the use of persistent homology for the identification of spatially dependent features in geographical data. In particular, they propose a one-sample permutation testing approach using persistence summary statistics [11]. Their finding suggest that persistent homology performs well at testing for spatial dependence in the presence of outliers or on sparsely connected graphs resembling ladder graphs [11]. Further, they conclude that there is potential for future work to investigate the use of alternative functional summaries of persistent homology. Addressing this outlook and widening the application to spatial omics data, we now investigate the use of persistence curves for SVG detection.
@@ 3 Background on Persistent Homology
3.1 Persistent Homology and Vertex-based Filtrations on Spatial Graphs
Throughout this paper, we set out to study a spatial dataset $X=\{x_{1},x_{2},...,x_{n}\}$ whose observations are assigned feature values $f:X\rightarrow\mathds{R}$ and spatial locations $s:X\rightarrow\mathds{R}^{2}$. We then want to examine the spatial distribution of such a feature, which is modelled by constructing a spatial (neighbourhood) graph $G$. This approach is appealing because adjacency relationships play a central role in spatial data analysis and we can flexibly construct a graph that best describes the nature of the data as also commonly done for analysing spatial omics data. See Section 4.4 for a more detailed description on our method.
@@ 3.1 Persistent Homology and Vertex-based Filtrations on Spatial Graphs
Persistent homology can then be computed from such a spatial graph by first considering the clique-complex $S=\operatorname{\mathrm{Cl}}(G)$ defined as the simplicial complex made up of cliques of vertices in $G$. Here, a clique complex is a generalisation of a graph that consists of 0-simplices corresponding to vertices, 1-simplices corresponding to edges, 2-simplices corresponding to 3-cliques (i.e. triangles), and so on. In particular, this simplicial complex is defined to be a set of simplices that is closed under taking intersections and face decomposition. In practice, we can consider $S$ as a generalised triangulation of a spatial graph.
@@
As a multi-scale summary of geometry, persistent homology then studies a filtration on this complex. That is, we construct a sequence of topological spaces, and track the evolution of topological features given by homology groups. The added algebraic structure allows for the computation of boundaries and holes from the generalised graph. We can thus track the emergence and disappearance of connected components, that is 0th dimensional homology groups, $H_{0}(S)$, and more generally k-dimensional homology groups, $H_{k}(S)$, representing $k$-dimensional holes [22]. This process as well as the further details explained below are illustrated via an example in Fig. 1.
@@
While there exist a multitude of methods for defining filtrations on graphs [3], we are interested in one specific construction throughout this work. In particular, we track the evolution of superlevel-sets defined from the spatial vertex feature, $f$, to investigate the spatial distribution of this feature. That is, we filter through the simplicial complex as follows
$S_{\delta}=Cl(G_{\delta})=\{s\in Cl(G)\mid f(u_{i})\geq\delta\text{ for all }u_{i}\in s\}$ (3.1)
@@
so that the filtration parameter $\delta$ decreases from max(f) to min(f). Here, $s$ is a simplex in the clique complex $S_{\delta}$ of the subgraph $G_{\delta}\in G$ defined as $G_{\delta}=\{u_{i}\in G\mid f(u_{i})\geq\delta\}$. The collection of simplicial complexes $S_{\max(f)}\subset...\subset S_{\min(f)}=Cl(G)$ then gives a Vietoris-Rips filtration. Intuitively, this vertex-based filtration can be constructed by filtering through the feature values starting with an empty simplicial set for $\delta>\max(f)$. Then, as $\delta$ decreases, a vertex $x_{i}$ i.e a zero-dimensional simplex is included whenever its associated vertex feature value $f(x_{i})$ has been reached. Further, an edge i.e. a one-dimensional simplex, is added to the filtered simplicial complex $S_{\delta}$ whenever both its vertices have been included [3, 33], with the process generalising to higher dimensions. Note that to define a sublevelset-set filtration, this construction can be modified, so that $\delta$ increases from $\min(f)$ to $\max(f)$. Further, for data collected on a rectangular grid, cubical persistence, which filters through cubical complexes made up of cubical cells, offers a natural alternative to the triangle-based simplicial complex construction. See Wagner et al. [41] for more details on how to generalise to this setting.
@@
The output from computing these filtrations results in a set of persistence diagrams, $\{\operatorname{\mathrm{PD}}_{0},\operatorname{\mathrm{PD}}_{1}\}$, one for each homology dimension, where each $\operatorname{\mathrm{PD}}_{\cdot}=\{(p_{1},q_{1}),(p_{2},q_{2}),...,(p_{m},q_{m})\}$ is a multi-set of persistence pairs in $\mathds{R}^{2}$ encoding the birth and death, i.e. the emergence and merging scales of topological features in the filtration. In particular, peaks or local maxima in spatial feature values are represented by points in the 0-th dimensional persistence diagram above the diagonal [22]. For a deeper exploration of the mathematical framework and theoretical underpinnings of persistent homology, we refer the interested reader to Edelsbrunner et al. [15]. We however move on to consider the use of persistence homology for statistical analyses.

### v2

@@ Abstract
Evaluating spatial patterns in data is an integral task across various domains, including geostatistics, astronomy, and spatial tissue biology. The analysis of spatial transcriptomics data in particular relies on methods for detecting spatially dependent features that exhibit significant spatial patterns for both explanatory analysis and feature selection. Given the complex and high-dimensional nature of these data, there is a need for robust, stable, and reliable descriptors of spatial dependence. We leverage the stability and multi-scale properties of persistent homology to address this task. To this end, we introduce a novel framework using functional topological summaries, such as Betti curves and persistence landscapes, for identifying and describing non-random patterns in spatial data. In particular, we propose a non-parametric one-sample permutation test for spatial dependence and investigate its utility across both simulated and real spatial omics data. Our vectorised approach outperforms baseline methods at accurately detecting spatial dependence. Further, we find that our method is more robust to outliers than alternative tests using Moran’s I.
@@ 1 Introduction
Omics data, that is measurements of molecules within an organism or a cell, provide crucial insights into biological processes and cellular function [45]. Specifically, transcriptomics plays a vital role in studying RNA transcripts, i.e. the expression of genes via the transcriptome. As a fundamental building block of molecular biology, transcriptomic measurements provide snapshots of cellular function and the mechanisms behind health and disease [28]. Recent advancements in spatial transcriptomics further enable the large-scale analysis of cells’ gene expression and their spatial locations across samples or within a tissue. Leveraging these data, the detection of spatially variable genes (SVGs) is an important analysis step. Spatially dependent transcripts are used both for explanatory analysis and for further downstream tasks, such as spatial domain identification and gene enrichment analysis [28, 20]. Thus, SVG detection is crucial for uncovering the spatial components of tissue biology. However, accurately identifying the spatial patterns in omics data on both local and global scales remains challenging, particularly due to the high-dimensional, sparse, and noisy nature of these data [5]. Challenges include the difficulty of modelling technical and systematic errors, the ongoing debate on even basic preprocessing steps, and an ever-growing understanding of the fundamental processes in cellular biology [20]. These hurdles necessitate the development of robust and expressive analysis methods [33]. In particular, SVG detection methods should be robust to technical noise and sparsity, correctly identify spatially dependent genes, be independent of the sum of expression values, and show low false-positive error rates [13, 25]. In practice, reliable SVG detection methods then help identify the spatial patterns of disease. For example, the analysis of spatially dependent transcripts enables the detection of spatial clusters of cancer cells or necrotising tissue within an organ. Further, SVG detection helps identify the genes driving these spatial changes [28].
Topological data analysis (TDA) has given scholars a new toolbox for quantifying complex structures in biological networks at multiple scales, motivated by the rigorous mathematical study of shapes [17]. Indeed, when adjusting for noise, molecular data in particular is equipped with meaningful coarse geometry [33]. TDA is thus poised to quantify spatial patterns in molecular data by capturing both local and global characteristics of the shape of data. Computing multi-scale topological features on spatial graphs gives stable and flexible summaries of the distribution of spatial features [11], summarising both local and global patterns in gene expression. In particular, our work demonstrates how topological descriptors can be used to distinguish spatial signals in transcriptomics data and identify genes whose expression varies across space. Our contributions are as follows:
@@ 2 Related Work
2.1 Spatially Variable Gene Detection
@@
The study of omics data underpins our understanding of fundamental biological processes [45]. The term ‘omics’ denotes molecular measurements in general, including features extracted from the genome, transcriptome, or proteome. As a key mechanism in biology, the set of DNA encoded in the genome is read out or transcribed by RNA molecules, which make up the transcriptome. Some of these transcripts later code for proteins, which are of great practical importance as they perform the actual biological functions in cells. Omics measurements then give key insights into key aspects such as disease pathology, cellular interactions, and disease development [28, 34]. In the last decade, rapid advancements in sequencing technologies have enabled the rapid growth in the collection and analysis of spatial omics data [27]. Spatial data are particularly interesting as they provide information on both molecular measurements as well as spatial tissue organisation. Observations from spatial omics datasets either correspond to individual cells, enabling the advent of single cell omics, tissue regions spanning multiple cells, or even sub-cellular measurements. Associated spatial locations are given by unstructured, i.e. continuous coordinates on a tissue sample, or spots on a pre-defined grid [28]. The focus of our investigation in particular are transcriptomic measurements that summarise the number of times a gene is measured by the number of RNA fragments detected at each capture location.
As a key processing step, transcriptomics data allows practitioners to detect genes whose expression varies notable between dissociated cells as the most highly variable cells of interest [20]. Spatial transcriptomic further allows us to determine genes whose expression varies in space. Such spatially variable genes then give key information on biological processes driving spatial tissue architecture, which motivates the development of a range of computational methods for detecting spatially variable genes [29, SVGs].
@@
Generally, SVG detection methods fall into numerous categories. Moran’s I is the most established statistical measure for spatial autocorrelation across spatial statistics [26]. As well as its cousin Geary’s C [18], Moran’s I is frequently used as a versatile descriptor of spatial patterns in gene expression values [20, 25]. Despite its simplicity as a mean summary based on weighted variance and covariance estimates, and its well-established theoretical benefits, Moran’s I is known to be sensitive to outliers and noise [43, 10], a disadvantage given the noisy nature of spatial omics data [13]. Further, based on the rich tradition of spatial statistics, there also exists a variety of statistical methods that fit spatial regression models to assess the importance of spatial covariance on gene expression. The most popular of these approaches include SpatialDE and SpatialDE2, which fit Gaussian regression processes, and SPARK or SPARK-X, fitting mixed and non-parametric models, respectively [14]. However, these models rely on fixed assumptions on the nature of the distribution of gene expression values, which remains a contested question in practice [47]. Lastly, there exist non-parametric methods, such as trendsceek, which models gene expression as a marked point process and uses permutation testing to make inferences [16]. This approach benefits from allowing statistical reasoning under less stringent assumptions but is associated with increased computational costs. Further, model-free frameworks, such as sepal [6], which measures diffusion times, or GLISS, which models graph Laplacians, have also been proposed [1]. However, despite the appeal of such geometric and model-free approaches, multi-scale descriptors, such as persistent homology, have not yet been employed in this context. Motivated by the use of geometric descriptors of spatial graphs, we thus aim to further develop persistent homology as a framework for spatial significance testing and focus on applications to spatial omics data.
@@ 2.2 Persistent Homology
Originating from algebraic topology, persistent homology (PH) enables practitioners to measure topological features from data [15], thus describing the shape of data [5]. This is of particular importance for molecular data, which, after accounting for noise, exhibits meaningful coarse geometry [33]. PH then provides one such a tool for encoding this geometry in a robust manner.
From a signal-processing perspective, persistent homology has demonstrated its utility for detecting peaks in noisy signals, guaranteeing stability properties and robustness to small perturbations [22, 21, 17]. The multi-scale nature of filtrations allows practitioners to summarise not only local but also global patterns in spatial data. Persistence-based methods can thus be applied to detect and encode both local and global extrema alike [21]. Indeed, the use of persistent homology as spatial descriptor has been extended to investigating time-varying data and spatial clustering [21, 30, 35]. Thus, persistence has shown its utility for describing spatial patterns, such as local maxima and local minima, in various contexts. For these reasons, persistence shows especially promising potential for SVG detection.
From a statistical perspective, the increasingly widespread use of PH for data analysis and budding applications to machine learning have also been driven by advancements in the statistical theory, underpinning the theoretical foundation of PH. In particular, summaries of PH allow for the definition of parametric and non-parametric significance tests and uncertainty estimation. Amongst these approaches, non-parametric techniques, specifically permutation testing, are by far the most utilised and assumption-free methods used for statistical inferences [9, 38]. Arguably, one-sample testing remains somewhat underexplored [11], with most common statistical approaches focusing on two-sample comparisons.
Nevertheless, recent work by Byers et al. [11] demonstrates the use of persistent homology for the identification of spatially dependent features in geographical data. In particular, they propose a one-sample permutation testing approach using persistence summary statistics [11]. Their findings suggest that persistent homology performs well at testing for spatial dependence in the presence of outliers or on sparsely connected graphs resembling ladder graphs [11]. Further, they conclude that there is potential for future work to investigate the use of alternative functional summaries of persistent homology. Addressing this outlook and widening the application to spatial omics data, we now investigate the use of persistence curves for SVG detection.
@@ 3 Background on Persistent Homology
3.1 Spatial Data as Spatial Graphs
Throughout this paper, we study a set of observations $X=\{x_{1},x_{2},...,x_{n}\}$, which correspond to cells or spatial regions on a tissue sample. Each observation is further associated with a spatial location $l(x_{i})\in\mathds{R}^{2}\text{ for }x_{i}\in X$. We are interested in assessing the spatial distribution of one gene or feature $f(x_{i})=f_{i}\in\mathds{R}\text{ for }x_{i}\in X$. This feature corresponds to measurements of one specific mRNA molecule of interest across a tissue slice. That is, the value $f_{i}$ measures the level of gene expression of one specific transcript at a spatial observation $x_{i}\in X$. We then want to examine the spatial distribution of this gene, which is modelled by constructing a spatial (neighbourhood) graph $G=(X,E)$, so that each vertex corresponds to an observation in $X$ and the edge set is constructed from the spatial locations. For example, an $\epsilon$-neighbourhood graph can be constructed as
$E=\{(x_{i},x_{j})\in X\times X\mid d(l(x_{i}),l(x_{j}))\leq\epsilon\},$ (1)
where $d:\mathds{R}^{2}\rightarrow\mathds{R}$ is a spatial distance metric between observations and $\epsilon\in\mathds{R}$ is some neighbourhood radius. This approach is appealing because adjacency relationships play a central role in spatial data analysis. By choosing a suitable neighbourhood criterion, we can flexibly construct a graph that best describes the nature of the data. See Section 4.4 for a more detailed description on our method. As an example illustrating our methodology, Fig. 1.1. shows a spatial neighbourhood graph, where the colour of each vertex indicates the expression values of one exemplary gene measured at that spatial location.
@@ 3.2 Persistent Homology and Vertex-based Filtrations
Persistent homology can be computed from a spatial graph $G=(X,E)$ by considering the clique-complex $S=\operatorname{\mathrm{Cl}}(G)$ defined as the simplicial complex made up of cliques of vertices. Here, a clique complex is a generalisation of a graph that consists of 0-simplices corresponding to vertices, 1-simplices corresponding to edges, 2-simplices corresponding to 3-cliques (i.e. triangles), and so on. More formally, a $k$-dimensional simplex in $S$ corresponds to a set of $k+1$ vertices that form a complete subgraph. Specifically, we have
$\operatorname{\mathrm{Cl}}(G)=\{\sigma\subseteq X\mid G[\sigma]\text{ is a complete graph}\},$ (2)
where $G[\sigma]$ is the subgraph of the graph $G$, induced by the vertex set $\sigma$. The clique complex defines a simplicial complex, i.e. a set of simplices that is closed under taking intersections and face decomposition, as any subset of a clique is also a clique itself. In practice, we can think of $S$ as a generalised triangulation of a spatial graph.
@@
As a multi-scale summary of geometry, persistent homology then studies a filtration on this complex. That is, we construct a sequence of topological spaces, and track the evolution of topological features given by homology groups. The added algebraic structure allows for the computation of boundaries and holes from the generalised graph. We can thus track the emergence and disappearance of connected components, that is zero-dimensional homology groups, $H_{0}(S)$, and more generally, $k$-dimensional homology groups, $H_{k}(S)$, representing $k$-dimensional holes [22]. This process, as well as the further details explained below, are illustrated via an example in Fig. 1.2. which plots the evolution of connected components (in blue) and one-dimensional loops (in pink) across a filtration on a spatial graph.
@@
While there exists a multitude of methods for defining filtrations on graphs [3], we are interested in one specific construction throughout this work. In particular, we track the evolution of superlevel-sets defined from the spatial vertex feature $f\colon X\rightarrow\mathds{R}$, as previously defined in Section 3.1. This filtration allows use to investigate the spatial distribution of one gene as we vary from the highest to the lowest expression value. That is, we filter through the simplicial complex as follows
$S_{\delta}=\operatorname{\mathrm{Cl}}(G_{\delta})=\{s\in Cl(G)\mid f(x_{i})\geq\delta\text{ for all }x_{i}\in s\}$ (3)
@@
so that the filtration parameter $\delta\in[\max(f_{i}),\min(f_{i})]$ decreases from the maximum to the minimum feature value. Here, $s$ is a simplex in the clique complex $S_{\delta}$ of the subgraph $G_{\delta}\in G$ defined as $G_{\delta}=G[\{x_{i}\in X\mid f(x_{i})\geq\delta\}]$. The collection of simplicial complexes $S_{\max(f_{i})}\subset...\subset S_{\min(f_{i})}=Cl(G)$ then gives a filtration of nested simplicial complexes. Intuitively, this vertex-based filtration can be constructed by filtering through the feature values starting with an empty simplicial set for $\delta>\max(f_{i})$. Then, as $\delta$ decreases, a vertex $x_{i}\in X$, i.e a zero-dimensional simplex, is included whenever its associated vertex feature value $f(x_{i})$ has been reached. Further, an edge, i.e. a one-dimensional simplex, is added to the filtered simplicial complex $S_{\delta}$ whenever both its vertices have been included [3, 36], with the process generalising to higher dimensions. Fig. 1.2. illustrates an example of a superlevel-set filtration constructed in this manner. Note that to define a sublevel-set filtration, this construction can be modified, so that $\delta$ increases from $\min(f_{i})$ to $\max(f_{i})$. Further, for data collected on a rectangular grid, cubical persistence, which filters through cubical complexes made up of cubical cells, offers a natural alternative to the triangle-based simplicial complex construction. See Wagner et al. [46] for more details on how to generalise to this setting.
@@
The output from computing these filtrations results in a set of persistence diagrams, $\{\operatorname{\mathcal{D}}_{0},\operatorname{\mathcal{D}}_{1}\}$, one for each homology dimension $k$, where each $\operatorname{\mathcal{D}}_{k}=\{(p_{1},q_{1}),(p_{2},q_{2}),...,(p_{m},q_{m})\}$ is a multi-set of persistence pairs in $\mathds{R}^{2}$ encoding the birth $p_{i}$ and death $q_{i}$, i.e. the emergence and merging scales of topological features in the filtration. In particular, peaks or local maxima in spatial feature values are represented by points in the 0-th dimensional persistence diagram above the diagonal [22]. This is illustrated in an example of a persistence diagram in Fig. 1.3., where the two persistence pairs plotted in blue correspond to the connected components that emerge during the filtration from the highest to the lowest gene expression values. The persistence diagram thus encodes spatial clustering behaviours in gene expression values which we want to leverage for detecting spatial dependence. For a deeper exploration of the mathematical framework and theoretical underpinnings of persistent homology, we refer the interested reader to Edelsbrunner et al. [15]. We, however, move on to consider the use of persistent homology for statistical analyses.